Using the Gram-Schmidt prodecure find an orthornornal set of vectors corresponding to the ordered basis B {(1,1,1),(1,1,0),(1,0,0)} of R^3. Also find a basis dual to B.


1
Expert's answer
2021-02-28T06:26:17-0500


Answer


a)

If { v1, v2, v3) is any ordered basis, then according to Gram-Schmidt procedure,


uk=vk- "\displaystyle\sum_{i=1}^{k-1}" proj uj (vk), where proju (v) = "u.v \over u.u"u.


and the normalized vector is ek= "u_k \over \sqrt{u_x.u_x }"


Here v1= "\begin{bmatrix}
 1 \\
 1 \\
1
\end{bmatrix}", v2"\begin{bmatrix}
 1 \\
 1 \\
0
\end{bmatrix}" v3"\begin{bmatrix}
 1 \\
 0 \\
0
\end{bmatrix}"



Now u1=v1"\begin{bmatrix}
 1 \\
 1 \\
1
\end{bmatrix}" and e1"u_1 \over \sqrt{u_1.u_1 }"= "1 \\over \\sqrt{3 }\n\n\n\u200b\n\n\n\n\u200b""\begin{bmatrix}
 1 \\
 1 \\
1
\end{bmatrix}""\begin{bmatrix}
 1 \\
 1 \\
1
\end{bmatrix}"




u2= v2- "u_1.v_2 \over u_1.u_1"u. = "\begin{bmatrix}
 1 \\
 1 \\
0
\end{bmatrix}" - "2 \over 3"    "\begin{bmatrix}
 1 \\
 1 \\
1
\end{bmatrix}" = "\begin{bmatrix}
 2 \over 3 \\
 2 \over 3 \\
2 \over 3
\end{bmatrix}"  



u2.u2= "1 \over 9""1 \over 9""4 \over 9""6 \over 9""2 \over 3"


e1= "\sqrt{3} \over \sqrt{2}""\begin{bmatrix}
 1 \over 3 \\
 1 \over 3 \\
-2 \over 3
\end{bmatrix}"


u3= v3- "u_1.v_3 \over u_1.u_1"u1 - "u_2.v_3 \over u_2.u_2"u2



u3"\begin{bmatrix}
 1 \\
 0 \\
0
\end{bmatrix}" - "1 \over 3" "\begin{bmatrix}
 1 \\
 1 \\
1
\end{bmatrix}" - "1 \over 2""\begin{bmatrix}
 1 \over 3 \\
 1 \over 3 \\
-2 \over 3
\end{bmatrix}"



u1.v3 = 1


u2.v3 = "1 \over 3"


u1.u1 = 3



u2.u2 = "2 \over 3"



b)


By definition, if B = { u1,u2, ---,uk,} is basis of a vector space, then

B* = { f1,f2, ---,fk,} is dual basis of dual space,


where fi (uj)=dkij , dkij= "\begin{cases}
 1 &\text{if } i=j \\
 0 &\text{if } i\not =j
\end{cases}"



Now given basis is B= {(1,1,1), (1,1,0), (1,0,0)}, then {  u1,u2,u3}


We have (u,y,z) = z (1,1,1)+ (y-z)(1,1,0) + (u-y)(1,0,0)


Now f1(u,y,z) =  z f1 (1,1,1)+ (y-z) f1(1,1,0) + (u-y) f1(1,0,0)

=z+ (y-z).0+ (u-y).0


So  f1(u,y,z) =  z


 f2(u,y,z) =  z  f2 (1,1,1)+ (y-z) f2(1,1,0) + (u-y) f2(1,0,0)

= z.0+ (y-z).1+ (u-y).0


so  f2(u,y,z) = y-z


 f3(u,y,z) =  z   f3 (1,1,1)+ (y-z) f3(1,1,0) + (u-y)  f3(1,0,0)

               = z.0+ (y-z).0+ (u-y).1


so  f3(u,y,z) = u-y


Thus  B* = { f1,f2, ,f3), where  f1(u,y,z) = z, f2(u,y,z) = y-z, f3(u,y,z) = u-y.